Grade 11Physical Chemistry

Calculate the ratio of degree of dissociation (α1/α2) when 1 M acetic acid solution is diluted to 1/100 times where Ka=1.8*10^-5

Assim Ujjwal
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
As we know
 
alpha = sqrt (K /C)
 
Hence
 
Alpha1 /alpha2 = sqrt(C 2 / C 1 ) = sqrt(100)
 
 = 10
 
Hope it helps
 
Regards
Arun